OpenStudy (anonymous):

In a 30°-60°-90° triangle, the length of the hypotenuse is 30. Find the length of the longer leg.

hartnn (hartnn):

longer leg = side opposite to 60 degree angle. use the property that side opposite to 60 degree angle = \(\large \dfrac{\sqrt 3}{2} \times hypotenuse\)